### v4

# Scalable Oversight for Superhuman AI via Recursive Self-Critiquing

Abstract
As AI capabilities increasingly surpass human proficiency in complex tasks, current alignment techniques, including SFT and RLHF, face fundamental challenges in ensuring reliable oversight. These methods rely on direct human assessment and become impractical when AI outputs exceed human cognitive thresholds. In response to this challenge, we explore two hypotheses: (1) Critique of critique can be easier than critique itself, extending the widely-accepted observation that verification is easier than generation to the critique domain, as critique itself is a specialized form of generation; (2) This difficulty relationship holds recursively, suggesting that when direct evaluation is infeasible, performing higher-order critiques (e.g., critique of critique of critique) offers a more tractable supervision pathway. We conduct Human-Human, Human-AI, and AI-AI experiments to investigate the potential of recursive self-critiquing for AI supervision. Our results highlight recursive critique as a promising approach for scalable AI oversight.
Machine Learning, ICML
1 Introduction
Supervision signals are fundamental to AI alignment (Bowman et al., 2022), providing the ground truth or preference data necessary to train models that behave in accordance with human expectations. The nature and accessibility of these supervision signals, however, vary substantially across different application domains. From a supervision acquisition perspective, tasks can be categorized into two types: (1) tasks with well-defined criteria, where ground truth can be deterministically obtained with low computational overhead, e.g., Go games and mathematical problems (Silver et al., 2017; Lightman et al., 2023); (2) tasks involving subjectivity or complex evaluation frameworks, such as business strategy and product design (Ouyang et al., 2022). The latter type is more prevalent in real-world applications and predominantly relies on human assessment, presenting a fundamental challenge.
Current alignment techniques, particularly Supervised Fine-tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF), have achieved empirical success with large language models (Meta, 2024; Yang et al., 2024; DeepSeek-AI, 2024). SFT (Chung et al., 2022; Wei et al., 2022) finetunes models with human-annotated demonstrations, showing particular efficacy in tasks where humans can effectively showcase desired behaviors. RLHF (Christiano et al., 2023; Ouyang et al., 2022) employs reinforcement learning with human preference reward models based on pairwise comparisons, extending supervision to more complex tasks where direct solution generation is challenging.
However, both approaches rely on direct human feedback, making them unsustainable for tasks where human evaluation becomes infeasible. For example, humans can struggle with time-consuming tasks such as reviewing extensive long-form text (Stiennon et al., 2022)r expertise-intensive tasks such as verifying solutions to complex mathematical problems (Li et al., 2024b). Furthermore, as AI capabilities advance beyond human abilities, obtaining reliable supervision signals becomes increasingly challenging, representing the central problem of scalable oversight (Casper et al., 2023; Ji et al., 2024; Kenton et al., 2024b).
The underlying insight of RLHF is that verification is easier than generation (Leike et al., 2018; Irving et al., 2018b). By recognizing critique as a specialized form of generation, we further hypothesize that critique of critique is easier than critique itself. Taking a complex mathematical proof as an example: while direct review can be challenging, assessing its critique is more manageable, as the key steps have already been identified. Moreover, we hypothesize that this difficulty relationship generalizes recursively, where each successive level of meta-evaluation becomes increasingly tractable. This resembles organizational decision-making processes, where managers evaluate their subordinates’ assessments rather than directly reviewing complex details. These hypotheses, if validated, offer a promising pathway for scalable oversight: while directly evaluating sophisticated AI output may exceed human capabilities, performing higher-order critiques could remain feasible.
To systematically verify these hypotheses, we first conduct Human-Human experiments where humans evaluate human outputs. We examine the progression from response to critique and then to critique-of-critique ($C^{2}$). By comparing accuracy under similar computational effort, completion time, and confidence levels, we find that higher-order critiques contribute to more effective evaluation than direct assessment. Furthermore, we demonstrate the recursive nature of this relationship by extending experiments to deeper critique chains, i.e., critique of critique of critique ($C^{3}$). Inspired by these human-human findings, we further investigate their applicability for supervising AI: when AI generates self-recursive critiques, can humans provide effective oversight by evaluating these critique chains? To answer this question, we conduct Human-AI experiments, where humans evaluate AI outputs on tasks where AI outperforms average humans. The results are promising across models of varying capabilities. Finally, we examine whether AI can achieve effective oversight through recursive self-critiques in AI-AI experiments across models of different capabilities. Our results demonstrate that recursive self-critiquing is effective in weak-to-strong scenarios, while the optimal critique strategy depends on the relative capabilities between supervised and critic models.
In general, our contributions can be summarized as follows:
  1. 1.
    We investigate and validate the hypothesis that critique of critique is easier than critique, extending the principle that verification is easier than generation.
  2. 2.
    We demonstrate that above difficulty relationship can hold recursively, showing how complex evaluation tasks can be simplified by recursive meta evaluations.
  3. 3.
    Through comprehensive Human-Human, Human-AI, and AI-AI experiments, we demonstrate the potential of recursive self-critiquing as a scalable oversight method, providing new valuable insights for supervising advanced AI systems beyond human capabilities.
2 Recursive Self-Critiquing
In this section, we introduce the protocols for recursive self-critiquing across multiple evaluation levels, spanning initial response through higher-order critiques. We then present majority voting and naive voting as two representative baselines to provide fair comparisons for evaluating the effectiveness of recursive critique.
Figure 1: Overview of the recursive critique framework. Starting from response generation for a given question, each subsequent level performs pair-wise evaluation of outputs from the previous level, forming a recursive critique chain. $C^{1}$ denotes Critique, $C^{2}$ denotes Critique of Critique, $C^{3}$ denotes Critique of Critique of Critique.
2.1 Protocols
As shown in Figure 1, the hierarchical criticism architecture progresses through multiple levels: from initial response, through first-order critique, to second-order critique of critique ($C^{2}$) and higher-order critiques. Our protocols follow standard RLHF practice (Ouyang et al., 2022), employing pairwise comparisons at each critique level. This approach leverages humans’ cognitive advantage in relative assessment over absolute evaluation (Jones and Inglis, 2015; Kelly et al., 2022), making recursive evaluation more tractable at each level. Moreover, this design facilitates consistency between human and AI experiments, as the latter requires pairwise preference data for reward model training.
Response
Response is the initial attempt to answer the question, serving as the foundation of the critique chain. Each response comprises a complete solution process and its corresponding answer:
$R(Q)\rightarrow(T^{0},A^{0})$ (1)
where $Q$ denotes the input question, $T^{0}$ represents the solution process which may include reasoning steps, justifications, and intermediate calculations, and $A^{0}$ is the final answer. Including the full solution process rather than merely the final answer enables critiques to better assess the correctness of each response by examining logical consistency, key step validity, and other aspects of the solution.
Critique
The first-order critique evaluates pairs of candidate responses for a given input question, conducting comparative analysis and providing reasoned judgment:
$C^{1}(Q,R_{1},R_{2})\rightarrow(T^{1},A^{1})$ (2)
where $R_{1}$ and $R_{2}$ denote two candidate responses, $T^{1}$ represents the critique rationale explaining which response is better and why, and $A^{1}$ is the final answer determined based on the critique analysis.
Critique of critique
The second-order critique evaluates pairs of first-order critiques, extending the evaluation to a higher level of abstraction:
$C^{2}(Q,R_{1},R_{2},C^{1}_{1},C^{1}_{2})\rightarrow(T^{2},A^{2})$ (3)
where $C^{1}_{1}$ and $C^{1}_{2}$ are two first-order critiques of the original responses, $T^{2}$ represents the analysis comparing the quality and validity of these critiques, and $A^{2}$ denotes the final answer determined by the superior critique.
Higher-order critiques
The $n$-th order critique continues this recursive process, leveraging assessments from all previous levels for evaluating pairs of (n1)(n-1)-th order critiques and reaching conclusions at this level:
$C^{n}(Q,R_{1},R_{2},C^{1}_{1},C^{1}_{2},\ldots,C^{n-1}_{1},C^{n-1}_{2})\rightarrow(T^{n},A^{n})$ (4)
where $C^{n-1}_{1}$ and $C^{n-1}_{2}$ are two (n1)(n-1)-th order critiques, $T^{n}$ represents the analysis comparing these critiques, and $A^{n}$ denotes the final answer derived from this comprehensive evaluation.
2.2 Baselines
We introduce two representative baseline strategies for rigorous comparison with recursive critique. The first is majority voting, which selects the most frequent answer from multiple evaluations. This baseline ensures fair comparison under equivalent computational effort. The second is naive voting, which performs direct aggregation of all available judgments from previous stages. This approach verifies whether recursive critique generates meaningful insights beyond simple consensus.
Majority voting
Since higher-order critiques are based on lower-order evaluation results, direct comparison between them would be unfair due to differing computational costs. To verify that the recursive structure achieves performance improvements by reducing supervision difficulty rather than merely benefiting from increased computational effort, we compare higher-order critiques with lower-order critiques under approximately equivalent computational effort. We achieve this through majority voting baselines (Wang et al., 2023) that aggregate multiple lower-order evaluations to match the computational cost of higher-order critiques.
Specifically, let $\epsilon(\cdot)$ denote the computational overhead for each evaluation. In AI experiments, this typically represents inference cost, while in human experiments, it’s more closely captured by annotation time spent on each evaluation task. As presented in Figure 1, the total computational effort $E(\cdot)$ for different-order recursive critiques $C^{1}$, $C^{2}$, and $C^{3}$ can be estimated as:
$E(C^{1})$ $=2\epsilon(R)+\epsilon(C^{1})\approx 3\epsilon(R)$ (5)
$E(C^{2})$ $=2\epsilon(R)+2\epsilon(C^{1})+\epsilon(C^{2})\approx 5\epsilon(R)$
$E(C^{3})$ $=2\epsilon(R)+2\epsilon(C^{1})+2\epsilon(C^{2})+\epsilon(C^{3})\approx 7\epsilon(R)$
We then define majority voting. For level $l$, given a set of $n$ evaluations, the majority voting result is:
$\text{Major}_{n}^{l}(\mathcal{A})=\underset{a}{\operatorname{argmax}}\sum_{i=1}^{n}\mathbbm{1}(A^{l}_{i}=a)$ (6)
where $A^{l}_{i}$ represents the judgment from the $i$-th evaluation at level $l$, and $\mathbbm{1}(\cdot)$ is the indicator function. This formula counts the occurrences of each possible answer among the $n$ evaluations and selects the most frequent one as the final result. In case of ties, where multiple answers have the same highest frequency, one is randomly selected. To enable fair comparison with recursive critique at level $l$ under approximately equivalent effort, we calculate $\text{Major}_{n}^{k}$ where $k<l$ and $n=E(C^{l})/E(C^{k})$. For example, $C^{3}$ should be compared with $\text{Major}_{3}^{2}$ (majority voting among three $C^{2}$ critiques) and $\text{Major}_{5}^{1}$ (majority voting among five $C^{1}$ critiques). Critically, majority voting aggregates independent evaluations without the structured pairwise comparison that defines recursive critique, allowing us to isolate whether improvements stem from the recursive structure versus computational scaling.
Naive voting baseline
A natural strategy for higher-order critique is to simply aggregate all judgments from previous stages through voting, adding no new analysis but merely following the consensus. The naive voting is defined:
$C^{1}_{\text{naive}}(R_{1},R_{2})$ $\rightarrow\text{Major}(\{A^{0}_{1},A^{0}_{2}\})$ (7)
$C^{2}_{\text{naive}}(C^{1}_{1},C^{1}_{2})$ $\rightarrow\text{Major}(\{A^{0}_{1},A^{0}_{2},A^{1}_{1},A^{1}_{2}\})$
$C^{3}_{\text{naive}}(C^{2}_{1},C^{2}_{2})$ $\rightarrow\text{Major}(\{A^{0}_{1},A^{0}_{2},A^{1}_{1},A^{1}_{2},A^{2}_{1},A^{2}_{2}\})$
We introduce this as a baseline to verify that proposed recursive critique outputs new insights rather than just follow simple vote aggregation results.
3 Is Recursive Critique Increasingly Easier?
In this section, we validate the hypothesis that critique of critique is easier than direct critique and examine whether this difficulty relationship holds recursively. We conduct experiments across diverse tasks with human annotators of similar abilities, and record their accuracy, completion time, and annotator confidence for analysis.
Table 1: Human experiment results across response, critique, and $C^{2}$ stages for five tasks. Bold numbers indicate best performance. Majority Voting@$E5$ represents voting results with computational effort equivalent to 5 times of response. Metrics include average accuracy, voting accuracy, naive voting, confidence (1-5), and completion time (minutes).
Dataset Stage Accuracy Majority Voting@$E5$ Naive Voting Confidence (1-5) Time (min)
CET-6 Response 49.11 55.80 3.074 18.36
Critique 58.13 60.78 49.22 3.253 17.03
$C^{2}$ 60.94 56.25 3.516 15.82
GAOKAO Math Response 66.29 81.81 3.201 14.58
Critique 82.50 86.61 66.41 3.863 14.62
$C^{2}$ 90.62 81.25 3.979 15.48
GAOKAO Chinese Response 71.56 79.69 3.822 17.81
Critique 78.65 84.38 64.84 4.026 13.91
$C^{2}$ 84.38 77.34 4.078 10.25
Figure Reasoning Response 65.00 78.12 3.888 16.74
Critique 75.00 77.08 65.62 4.213 16.01
$C^{2}$ 79.69 72.66 4.313 15.02
KAOGONG Response 69.69 83.59 3.828 16.26
Critique 84.38 84.90 70.31 4.031 15.48
$C^{2}$ 85.94 82.81 4.031 12.58
3.1 Tasks
We select five representative tasks that require diverse cognitive capabilities while maintaining moderate difficulty. These tasks span multiple domains, including language comprehension, mathematical reasoning, logical analysis, and visual reasoning, to test the generalizability of recursive critique framework across different cognitive skills. All tasks include 64 multiple-choice questions. Each task consists of 64 multiple-choice questions.
CET-6
The College English Test Band 6 (CET-6) is a standardized English proficiency assessment for Chinese university students. From its Careful Reading section, we select one question per passage; each passage contains 400-450 words and includes multiple-choice questions that test main idea comprehension, vocabulary understanding, and inference abilities. This task requires English language proficiency, reading comprehension skills, and analytical reasoning to extract meaning from complex texts. As few of our annotators have passed CET-6, these questions present substantial challenges.
GAOKAO Chinese
The Chinese reading comprehension questions are drawn from China’s National College Entrance Examination (Gaokao). These questions demand accurate comprehension of the original text and logical reasoning capabilities for answer selection. Since our annotators are college graduates who previously took the Gaokao, these questions present moderate difficulty.
GAOKAO Math
The mathematics questions are sourced from standardized high school tests (Zhang et al., 2023). Since problem difficulty typically increases with question number and considering that our annotators graduated several years ago with some having non-science backgrounds, we select the first ten multiple-choice problems to ensure moderate difficulty for them. These questions require mastery of mathematical concepts and formulas as well as the ability to apply mathematical reasoning to solve problems.
KAOGONG
The questions are sourced from China’s National Civil Service Exam, the annual government recruitment test. These questions assess logical reasoning, language understanding, and numerical analysis skills. We exclude knowledge-based questions to focus on cognitive abilities requiring analytical thinking and problem-solving rather than factual recall.
Figure Reasoning
These visual tasks from the Civil Service Examination assess logical abilities through non-verbal reasoning without requiring domain-specific knowledge or cultural context, demanding spatial reasoning skills, pattern recognition, and abstract thinking capabilities.
3.2 Setup
Participants
We recruit 32 participants with bachelor’s degrees: 22 from STEM backgrounds and 10 from liberal arts backgrounds. Most participants passed CET-4 and scored approximately 100 out of 150 on their high school mathematics exams. All participants have prior experience in data annotation and are employed full-time for this study.
Execution
We develop standardized guidelines for all tasks using instructions and examples, detailed in Appendix A. Tasks are organized into data packages with specified submission deadlines, and annotators are randomly assigned across different critique levels to ensure participation in all stages. To maintain efficiency, we set a 20-minute time limit for each question at every stage, managed through flexible package-level deadlines that allow annotators to allocate time as needed. Annotators complete a predetermined number of tasks daily within their scheduled working hours. We conduct regular feedback sessions to gather suggestions for refining procedures and guidelines, and assign dedicated staff to oversee process management and quality control.
Metrics
We assess the effectiveness of recursive critique through three metrics: (1) accuracy measures consistency with ground truth answers; (2) completion time records the duration of the entire evaluation process; (3) confidence reflects participants’ self-assessed certainty in their final answers on a five-point scale.
3.3 Critique of Critique can be Easier than Critique
We validate the hypothesis that critique of critique is easier than critique across five tasks. The results in Table 1 show consistent improvements from response to critique to $C^{2}$ stages. Taking GAOKAO Math as an example, average accuracy improves from 66.29% (response) to 82.50% (critique) and further to 90.62% ($C^{2}$), while completion time remains stable or slightly decreases (e.g., from 18.36 to 15.82 minutes for CET-6). Under comparable effort, majority voting shows similar trends. For instance, accuracy improves from 81.81% (response) through 86.61% (critique) to 90.62% ($C^{2}$) in GAOKAO Math, demonstrating the advantage of higher-order critique. Compared to naive voting, average accuracy consistently outperforms. Taking GAOKAO Math as an example, naive voting achieves only 66.41% at the critique stage and 81.25% at $C^{2}$, significantly lower than the average accuracy of 90.62%. These results validate that recursive critique generates new insights rather than merely aggregating previous judgments. Moreover, annotator confidence shows steady improvement across stages, suggesting that higher-order critique becomes more tractable.
3.4 Recursive Critique Remains Consistently Easier
We extend recursive critique to third-order critique ($C^{3}$) on two representative tasks. As shown in Table 2, accuracy improves continuously at the $C^{3}$ level in both tasks, with CET-6 increasing from 60.94% at $C^{2}$ to 67.19%, and GAOKAO Math from 90.62% to 93.75%. Under comparable computational effort, majority voting shows similar improvements, reaching 67.19% for CET-6 and 93.75% for GAOKAO Math. However, naive voting achieves substantially lower performance than the average accuracy at each critique level. Meanwhile, confidence scores improve while completion time decreases across critique levels. These results demonstrate that recursive critique remains consistently easier and that its benefits extend beyond mere computational scaling or consensus aggregation.
Table 2: Human experiment results across response, critique, $C^{2}$, and $C^{3}$ stages for two tasks. Bold numbers indicate best performance. Majority Voting@$E7$ represents voting results with computational effort equivalent to 7 times of response. Metrics include accuracy, majority voting accuracy, naive voting, confidence (1-5), and completion time (minutes).
Dataset Stage Accuracy Majority Voting@$E7$ Naive Voting Confidence (1-5) Time (min)
CET-6 Response 49.11 57.03 3.074 18.35
Critique 58.13 63.28 49.22 3.253 17.03
$C^{2}$ 60.94 63.02 56.25 3.516 15.82
$C^{3}$ 67.19 60.16 3.766 14.23
GAOKAO Math Response 66.29 85.94 3.194 14.58
Critique 82.50 88.28 66.41 3.863 14.62
$C^{2}$ 90.62 91.15 81.25 3.979 15.48
$C^{3}$ 93.75 87.50 4.031 14.14
4 Can Recursive Self-Critiquing Enable Human Oversight of AI?
In this section, we further conduct Human-AI experiments to examine whether recursive critique enables effective human oversight when capabilities exceed human performance.
4.1 Tasks
We select tasks based on the criterion that humans find them challenging while AI demonstrates reasonable but not perfect performance, creating suitable conditions for meaningful evaluation of human oversight when AI capabilities exceed human performance. Following this criterion, we select two challenging task types for our experiments:
  • GAOKAO Math comprises the last two multiple-choice questions from the high school mathematics examination (Zhang et al., 2023), which demand advanced problem-solving skills and mathematical reasoning abilities.
  • TEM4 (Test for English Majors Grade Four) includes reading comprehension questions that require professional-level English proficiency and complex text analysis capabilities.
Both tasks are beyond most annotators’ abilities while remaining moderately challenging for AI models. We filter out questions where models achieve either 0% or 100% accuracy, as these overly easy or difficult cases produce uniform outputs that are unsuitable for our validation.
4.2 Setup
We employ the same annotators, annotation procedures, and evaluation metrics as in Human-Human experiments. The annotation process follows the Human-Human procedure, with AI outputs replacing human ones. To obtain AI responses, we utilize both Qwen-7B-Instruct and Qwen-72B-Instruct models (Qwen et al., 2025) to examine recursive critique across different AI capability levels. For each question, the AI model first generates initial responses, then performs self-critique recursively at multiple orders ($C^{1}$, $C^{2}$). Human annotators evaluate AI outputs at each corresponding stage, except for the Response stage where humans complete tasks independently without relying on AI outputs.
4.3 Experimental Results
Table 3: Performance comparison across recursive critique stages, with human accuracy subscripts showing difference from previous-stage AI accuracy. Results from Qwen2.5-7B/72B-Instruct on mathematics and English tests, including accuracy, confidence (1-5), and completion time (minutes).
Dataset Stage Human Accuracy AI Accuracy Confidence (1-5) Time (min)
GAOKAO Math (Qwen2.5-7B) Response 43.75 46.09 2.188 23.23
Critique 53.12+7.03 47.66 2.578 22.92
$C^{2}$ 56.25+8.59 50.78 3.156 23.91
$C^{3}$ 54.69+3.91 3.109 16.56
GAOKAO Math (Qwen2.5-72B) Response 43.75 63.28 2.188 23.23
Critique 68.75+5.47 61.72 3.375 25.41
$C^{2}$ 70.31+8.59 64.06 3.625 21.30
$C^{3}$ 65.62+1.56 3.469 22.94
TEM4 (Qwen2.5-7B) Response 34.38 52.34 3.234 22.44
Critique 59.38+7.04 61.72 3.750 17.55
$C^{2}$ 67.19+5.47 64.84 3.766 18.14
$C^{3}$ 64.06-0.78 3.797 16.52
TEM4 (Qwen2.5-72B) Response 34.38 65.62 3.234 22.44
Critique 67.19+1.57 65.62 3.875 16.56
$C^{2}$ 64.06-1.56 67.97 3.859 15.47
$C^{3}$ 71.88+3.91 3.813 16.86
Recursive critique enables effective human oversight of AI.
Table 3 indicates that human response accuracy is lower than AI accuracy, showing that AI surpasses humans in these tasks. However, in subsequent critique stages, humans consistently achieve higher accuracy than AI’s previous outputs. For example, with Qwen2.5-7B on GAOKAO Math, human accuracy reaches 53.12% at the critique stage (7.03% higher than AI’s initial 46.09%), and further increases to 56.25% at $C^{2}$ (8.59% above AI’s critique stage). This finding suggests that recursive critique enables human supervision even in the tasks which AI outperforms humans.
Recursive critique improves evaluation efficiency and confidence.
Despite processing more information at higher levels, completion time generally decreases or remains stable. For TEM4 with Qwen-72B, time decreases from 22.44 minutes at the response level to 15.47 minutes at $C^{2}$. Meanwhile, annotator confidence shows consistent improvement across levels and model scales, particularly in the mathematics task with Qwen-72B where confidence increases from 2.19 to 3.63. These results suggest that recursive critique makes evaluation more tractable.
5 Can Recursive Self-Critiquing Achieve Better AI Supervision?
In this section, we conduct AI-AI experiments to explore the potential of recursive self-critiquing for achieving better AI supervision under weak-to-strong, strong-to-weak, and self-supervised settings.
5.1 Setup
Model Preparation
We investigate the dynamics of supervisory effectiveness across pairings of supervised and critic models with different capability levels. We use models from the Qwen2.5 series (Qwen et al., 2025), operating under the premise that model capability generally correlates with parameter size. However, different variants of the Qwen2.5-Instruct series may have undergone different post-training procedures, which could introduce confounding factors. To ensure fair comparison, we randomly sample 282K instances from the open-source TULU-3-SFT dataset (Lambert et al., 2024) and fine-tune all models from the Qwen2.5-Base series using the same data.
Data Preperation
To ensure objective measurement of supervision quality, we select mathematical tasks due to their verifiable nature. The experimental data are drawn from the DeepScaleR dataset (Luo et al., 2025), with 512 randomly sampled instances as the test set and the remainder as training data. We employ the Math-Verify library (Kydlíček and Gandenberger, 2025) to determine answer correctness and obtain reliable ground truth signals.
Experiment Setting
In our experiments, the supervised model first performs recursive self-critique at varying orders. Subsequently, a critic model conducts a final higher-order critique based on the supervised model’s outputs. Prompts and sampling strategies are detailed in Appendix B. Following established RLHF methodologies (Ouyang et al., 2022), we use these final critiques to construct preference pairs and train reward models. To avoid potential confounding effects from architectural similarities, we select Llama3.1-8B (Meta, 2024)—a different architecture with comparable model capacity—as the foundation for the reward model. This reward model is then used for Best-of-N sampling to systematically evaluate supervisory efficacy across diverse model-critic combinations.
Evaluation Metric
To quantify supervision effectiveness, we adopt the Performance Recovered (PR) metric following the framework established by (Burns et al., 2023):
$\text{PR}=\frac{\mathbb{E}_{x\sim\mathcal{D}}[r^{*}(x,\arg\max_{y\in\{y_{i}\}_{i=1}^{n}}r(x,y))]}{{\mathbb{E}_{x\sim\mathcal{D}}[\max_{y\in\{y_{i}\}_{i=1}^{n}}r^{*}(x,y)]}}$ (8)
In this formulation, $x\sim\mathcal{D}$ denotes inputs sampled from distribution $\mathcal{D}$, and $\{y_{i}\}_{i=1}^{n}\sim M(\cdot|x)$ represents $n$ samples generated by model $M$ given input $x$. The learned reward function is denoted as $r(x,y)$, with $r^{*}(x,y)$ representing the ground truth reward function. Intuitively, the numerator measures the ground truth quality of outputs selected by the learned reward model, while the denominator represents the oracle performance with perfect selection. For mathematical tasks where $r^{*}$ indicates binary correctness, PR thus quantifies how close the learned reward model comes to oracle pass@N performance.
Figure 2: PR scores with a fixed 7B critic model and supervised models of varying sizes.
Figure 3: PR scores with a fixed 7B supervised model and critic models of varying sizes.
5.2 Experimental Results
Figures 3 and 3 present our experimental results under two settings: (1) Figure 3 shows results where supervised models of varying sizes first perform recursive self-critique, followed by evaluation from a fixed 7B critic model at each stage. The critic’s judgments are used to train reward models specific to each supervised model size, which then guide Best-of-N sampling on the corresponding supervised models. The PR metric compares this Best-of-N performance against the oracle Pass@N performance for each model size. (2) Figure 3 shows results where a fixed 7B supervised model first performs recursive self-critique, followed by evaluation from critic models of varying sizes at each stage. The critics’ judgments are used to train reward models specific to each critic size, which then guide Best-of-N sampling on the same 7B supervised model. The PR metric compares this Best-of-N performance against the oracle Pass@N of this fixed 7B model across different critic sizes.
Recursive self-critiquing benefits weak-to-strong supervision.
Figure 3 demonstrates that when supervised models are larger than the 7B critic model, higher-order critiques generally yield improved performance compared to lower-order critiques. Similarly, Figure 3 shows that when critic models are smaller than the 7B supervised model, higher-order recursive critiques can provide better supervision effectiveness. Both findings support recursive self-critiquing as a promising approach to scalable oversight, particularly in scenarios where humans (as the "weaker model") oversee increasingly capable AI systems (the stronger model).
Direct supervision exhibits superior performance in strong-to-weak settings.
Conversely, both 3 and 3 show that when critic models are stronger than the supervised model, direct critique produces better results than allowing the supervised model to engage in higher-order self-critique first. This asymmetry indicates that self-critique from weaker models may not be effective and can even mislead stronger critics.
6 Discussion and Related Work
Limitations in Current Alignment Strategies.
RLHF has emerged as the dominant approach in AI alignment, building upon the principle that "verification is easier than generation" (Irving et al., 2018b). However, it utilizes static reward models as proxies for human preferences and thus introduces the risk of reward hacking (Gao et al., 2022; Karwowski et al., 2023). While approaches such as iterative annotation and tool augmentation (Li et al., 2024a; Gou et al., 2024) provide intermediate solutions, they face limitations in supervision capability.
Related Approaches to Scalable Oversight.
As AI capabilities exceed human expertise, humans may no longer provide effective supervision (Amodei et al., 2016). One approach to this challenge is the debate protocol (Irving et al., 2018a), where agents argue for opposing answers (Khan et al., 2024; Michael et al., 2023), though with demonstrated limitations (Kenton et al., 2024a). Unlike debate’s adversarial framework, our approach assumes higher-order critic tasks are inherently easier. Task decomposition methods (Wu et al., 2021) break complex oversight into manageable sub-problems through breadth-first strategies, while our method employs depth-first recursive refinement. Our majority vote baseline builds on self-consistency methods (Wang et al., 2023; Fluri et al., 2023).
Mechanisms of Recursive Self-Critiquing and Implications.
The effectiveness of recursive self-critiquing stems from several mechanisms: higher-order criticism shifts attention from details to abstract principles, each critique level provides structured context, and the recursive structure transforms absolute tasks into pairwise judgments, leveraging humans’ advantage in relative assessment (Jones and Inglis, 2015; Kelly et al., 2022). However, our AI experiments show that direct supervision can be superior in strong-to-weak settings, which may be due to the limited critique capabilities of current models (Xi et al., 2024). Future work may focus on enhancing model critique capabilities (Wang et al., 2024a; Yu et al., 2025; Ankner et al., 2024).
7 Conclusion
This work investigates how to obtain reliable supervision signals when AI capabilities surpass human abilities. Through Human-Human, Human-AI, and AI-AI experiments, we examine the hypothesis that critique of critique is easier than critique and demonstrate that this difficulty relation holds recursively. The experiments suggest a promising pathway for scalable oversight through recursive self-critiquing when direct human evaluation becomes infeasible.
Impact Statement
Our recursive self-critiquing framework offers potential for maintaining effective AI oversight as capabilities surpass human abilities. However, this approach carries risks, including false confidence in oversight effectiveness, vulnerability to adversarial examples. Our experiments also reveal current limitations in AI models’ recursive self-critiquing capabilities, highlighting the need for continued development of models’ self-critique abilities to enhance oversight robustness. We acknowledge these potential impacts and encourage continued research to strengthen scalable oversight methods.
Appendix A Human Experiments Guidelines
This section details the guidelines and quality assurance involved in the Human-Human and Human-AI experiments. We establish consistent and comprehensive guidelines for annotation tasks at different stages across various tasks. Our guidelines emphasize the quality of the reasoning process over accuracy rates, requiring annotators to articulate their thinking process clearly without accessing external references. While accuracy is encouraged, the primary focus is on providing clear, well-reasoned justifications for their decisions. Annotators are instructed to invest their time primarily in analytical thinking, expressing their reasoning in clear, concise, and logically coherent natural language. The guidelines provide suggested formats but maintain flexibility, prioritizing the clear documentation of thought processes over rigid adherence to specific forms [^1] Fixed templates were initially tested but abandoned as annotators reported them to be unflexible and burdensome.. We provide detailed instructions at each stage in the following sections.
A.1 Response Stage
In the response stage, annotators are presented with a source text, a question, and multiple choice options. The primary task is to select the correct answer and provide comprehensive reasoning for their choice.
Recommmanded Annotation Template
The response should clearly indicate the selected answer and provide a complete reasoning process. This process should include specific citations from the source text as evidence, logical analysis that connects the evidence to the conclusion, and step-by-step reasoning where applicable. For example, responses can follow two primary patterns:
  • Option B is correct because [evidence + reasoning].
  • Options A/C/D are incorrect because [evidence + reasoning], therefore B is selected.
Other patterns are also acceptable as long as they maintain clear reasoning and sufficient evidence support. The examples of high-quality and low-quality responses are provided in Table 6 for illustration.
Quality Requirements
Response annotations must satisfy four fundamental criteria:
  • Relevance: Direct connection to the question and source text
  • Organization: Clear logical structure and information flow
  • Clarity: Concise expression without unnecessary complexity
  • Coherence: Smooth transitions between reasoning steps
A.2 Critique Stage Annotation
In the critique stage, annotators evaluate two responses from the previous stage based on the source text and question. The evaluation should focus on the correctness of responses, examining their logical coherence and evidence support.
Recommended Annotation Template
The critiques should clearly present the final judgment and supporting rationale with referenced evidence cited in the responses or the question. For example, common annotation patterns include:
  • Agreement with Response 1 with specific justification, noting uncertainties or disagreements with Response 2.
  • Agreement with Response 1 with justification, identifying specific errors in Response 2.
  • Agreement with both responses, providing supporting evidence for the shared conclusion.
  • Disagreement with both responses, detailing specific errors and providing justification for an alternative answer.
Critiques should prioritize identifying key errors that affect the final judgment, while minor issues that do not impact the conclusion are optional. The high quality and low quality examples is presented in Table 7 and Table 8.
Quality Requirements
critique annotations must satisfy five fundamental criteria:
  • Relevance: Direct connection to the question and source text
  • Organization: Clear logical structure and information flow
  • Clarity: Concise expression without unnecessary complexity
  • Coherence: Smooth transitions between reasoning steps
  • Objectivity: Fair analysis of responses’ strengths and weaknesses
A.3 Higher-Order Critique Stage
In the higher-order critique stage, annotators evaluate two critique annotations based on the source text, question, and responses. The evaluation should focus on assessing the critiques’ reasoning process, examining the validity of their evidence analysis, and identifying any logical gaps or oversights.
Recommended Annotation Template
The higher-order critiques should clearly present their evaluation of both critiques’ analyses and provide a final judgment with supporting rationale. For example, common annotation patterns include:
  • Agreement with Critic 1 with specific justification, noting uncertainties or disagreements with Critic 2.
  • Agreement with Critic 1 with justification, identifying specific errors in Critic 2’s analysis.
  • Agreement with both critics, acknowledging their shared valid points while noting potential weaknesses.
  • Disagreement with both critics, detailing specific logical flaws and providing independent justification.
Critics should prioritize identifying key errors in the critics’ reasoning while noting potential improvements even when agreeing with their conclusions.
Quality Requirements
Higher-order critique annotations must satisfy six fundamental criteria:
  • Relevance: Direct connection to the question and critics’ analyses.
  • Organization: Clear logical structure and information flow.
  • Clarity: Concise expression without unnecessary complexity.
  • Coherence: Smooth transitions between reasoning steps.
  • Objectivity: Fair analysis of critics’ strengths and weaknesses.
  • Improvement: Identification of gaps or potential enhancements in critics’ reasoning.
Examples of high-quality and low-quality higher-order critiques are presented in Tables 9 and 10.
Appendix B Prompts for AI-AI Experiments
We adopt the following prompt template in Figure 5, 6, 7, 8 to conduct response generation and multi-stage critiques. Additionally, our smaller SFT models, particularly those with 0.5B parameters and limited capabilities, occasionally fail to follow instructions properly. To address this issue, we incorporate hints in the output section to enhance the model’s instruction adherence and chain-of-thought analysis process. We set the sampling temperature to 1.0 and top_p to 1.0.
Appendix C Recursive Self-critiquing on Differernt LLMs
In this section, we further investigate the effectiveness of recursive self-critiquing across different LLMs on various tasks.
Table 4: Performance comparison of AI self recursive critiquing. We select the question set that $Q^{\prime}=\{q\mid 0<\text{Acc}(q)<0.7,q\in Q\}$ to focus on questions where initial accuracy is moderate, as questions with very high initial accuracy leave limited room for meaningful improvement through recursive self-critiquing.
Dataset Stage Qwen2.5-14B-Instruct Gemma2-9B-Instruct Qwen3-30B-A3B-Thinking-2507
Accuracy Majority Naive Accuracy Majority Naive Accuracy Majority Naive
MMLU Pro Response 34.71 35.58 22.31 25.43 37.21 36.80
Critic 35.50 35.58 35.17 32.95 32.81 28.90 41.62 41.98 40.00
$C^{2}$ 35.78 35.67 35.42 32.25 32.24 30.35 40.81 41.09 40.91
$C^{3}$ 36.83 35.25 31.79 31.04 37.15 37.56
BoolQ Response 25.98 20.41 31.36 28.78 33.40 28.09
Critic 27.14 24.49 27.35 32.59 30.24 31.22 33.70 32.68 32.15
$C^{2}$ 26.53 26.12 25.92 29.67 28.05 27.80 32.20 32.06 32.15
$C^{3}$ 28.16 25.51 32.44 27.07 34.03 32.87
MATH Response 31.69 31.19 22.82 19.90 52.56 61.86
Critic 34.56 34.81 34.27 26.14 25.23 23.30 50.85 52.27 52.65
$C^{2}$ 35.19 34.92 34.86 26.90 26.60 25.00 41.76 42.61 43.18
$C^{3}$ 35.89 35.41 27.32 25.69 36.08 36.36
GPQA Response 22.09 19.56 19.68 16.24 38.14 40.89
Critic 23.84 23.46 23.05 24.43 23.92 19.57 38.82 39.88 40.44
$C^{2}$ 23.30 23.24 22.50 22.60 22.31 20.39 37.74 37.55 37.34
$C^{3}$ 24.26 23.35 22.63 20.75 35.70 35.68
TruthfulQA Response 25.73 22.37 24.74 22.63 36.21 36.26
Critic 39.57 38.45 34.12 39.98 39.37 29.68 43.97 44.06 43.30
$C^{2}$ 37.87 37.84 34.54 34.67 35.68 30.74 44.18 44.25 43.97
$C^{3}$ 38.66 36.49 37.26 32.11 44.47 44.25
C.1 Datasets
We utilize reasoning, knowledge, and alignment-related datasets, including the following:
  • MATH(Hendrycks et al., 2021) is a mathematical problem-solving dataset consisting of 12,500 challenging competition-level math problems, designed to assess machine learning models’ mathematical reasoning abilities. Each problem is accompanied by a fully worked-out step-by-step solution, enabling models to learn how to generate answer derivations and explanations.
  • GPQA(Rein et al., 2023) is a highly challenging multiple-choice question dataset consisting of 448 questions crafted by domain experts in biology, physics, and chemistry. The dataset is designed to assess the reasoning capabilities of both human experts and state-of-the-art AI models on complex scientific topics. To ensure its difficulty and quality, questions were validated by experts with PhD-level knowledge, achieving an accuracy of only 65% (or 74% after correcting clear retrospective mistakes). In contrast, highly skilled non-expert validators, even with unrestricted web access for over 30 minutes per question, achieved only 34% accuracy.
  • TruthfulQA(Lin et al., 2022) evaluates the truthfulness of language models in answering questions, comprising 817 questions across 38 categories, including health, law, finance, and politics. The questions were carefully designed to reflect common human misconceptions or false beliefs, making them particularly challenging. To perform well, models must avoid generating false answers learned from imitating human-written text, which often contains misinformation.
  • BoolQ(Clark et al., 2019) is a reading comprehension dataset designed to study naturally occurring yes/no questions, meaning questions that arise spontaneously in unprompted and unconstrained settings. The dataset presents unexpected challenges, as its questions often involve complex, non-factoid information and require entailment-like inference rather than simple fact retrieval.
  • MMLU-Pro(Wang et al., 2024b) is an enhanced version of MMLU designed to go beyond MMLU’s primarily knowledge-driven evaluation. MMLU-Pro incorporates more challenging reasoning-focused questions, expands the answer choice set from 4 to 10 options, and removes trivial and noisy questions from MMLU. Experimental results show that MMLU-Pro significantly increases difficulty, leading to an accuracy drop of 16% to 33% compared to MMLU.
C.2 Setup
Models and Configuration
We conduct experiments with three state-of-the-art language models: Qwen2.5-14B-Instruct (Qwen et al., 2025), Gemma2-9B-Instruct (Team et al., ), and Qwen3-30B-A3B-Thinking-2507 (Team, 2025). For Qwen2.5-14B-Instruct and Gemma2-9B-Instruct, we set temperature=1.0 and top-p=0.95 to enhance sampling diversity. For Qwen3-30B-A3B-Thinking-2507, we follow official recommended settings with temperature=0.6, top-p=0.95, top-k=20, and min-p=0, extracting only the response content while excluding the thinking traces. We employ structured prompts to ensure consistent response formats across models and datasets (Figures 9, 10, 11, 12, 13, and 14), with minor model-specific adjustments for output length constraints and answer formatting.
Sampling Strategy and Evaluation
To ensure fair comparison across recursive critique stages, we follow a controlled hierarchical sampling strategy. For Qwen2.5-14B-Instruct and Gemma2-9B-Instruct, we generate 7 initial responses per question, derive 5 critics from the first two responses, then produce 3 critics of critics ($C^{2}$) and 1 critic of critics of critics ($C^{3}$). This process repeats 10 times per question with averaged results. For Qwen3-30B-A3B-Thinking-2507, due to longer output length and resource constraints, we use 16 initial responses with 4 critic pairs each producing 4 critiques at each subsequent level ($C^{2}$ and $C^{3}$). This pyramidal structure balances computational efficiency with evaluation diversity. We adopt identical evaluation metrics and baselines as used in the human experiments.
Table 5: Performance comparison by initial response correctness. C denotes correct, W denotes wrong responses. 1C1W/2C/2W indicate different combinations of initial response correctness. AI results from MATH dataset, human results from multiple datasets.
(a) AI performance
Stage Type Gemma2 Qwen2.5
Critic 1C1W 42.3 55.5
2C 64.3 98.4
2W 13.6 1.1
$C^{2}$ 1C1W 46.5 55.7
2C 89.8 97.1
2W 4.8 1.6
$C^{3}$ 1C1W 51.1 52.3
2C 92.8 98.9
2W 2.7 1.3
(b) Human performance
Stage Type Accuracy
Critic 1C1W 56.6
2C 86.8
2W 21.0
$C^{2}$ 1C1W 73.7
2C 93.8
2W 29.0
$C^{3}$ 1C1W 75.9
2C 93.5
2W 31.2
C.3 Experimental Results
Potential effectiveness in specific models.
The results in Table 4 compare the performance of Qwen and Gemma models across different datasets. From these results, we observe clear disparities in higher-order critiquing abilities across both models and benchmarks. Qwen2.5-14B-Instruct demonstrates stronger recursive-critiquing effectiveness, exhibiting consistent improvements from the initial response to deeper critique stages. However, other models, such as Gemma2-9B-Instruct or Qwen3-30B-A3B-Thinking-2507, show gains only on a limited subset of datasets.
These performance gaps likely stem from difficulties in distinguishing true statements from inputs containing mixed true–false information, as illustrated in Table 5. For comparison, Table 5 reports the accuracy of our human annotators across different input conditions. Human performance not only far exceeds that of AI models in both the 1C1W and 2W scenarios, but also shows substantial improvement as the length of the recursive-critique chain increases.
Current AI models show limited capabiliy in self-recursive critique.
We further investigate recursive self-critique performance across different large models and accuracy intervals. Testing models ranging from Qwen2.5-7B to 72B, we find that models typically demonstrate self-critique effectiveness in intervals where response accuracy is relatively moderate. However, overall we observe that models’ self-critique capabilities are limited, with typically modest improvement margins. These results are also partially validated in prior work (Huang et al., 2023; Tang et al., 2025) and summarized by Kamoi et al. (2024). This finding further highlights the importance of investigating approaches to improve models’ critique performance (McAleese et al., 2024).
Figure 4: The relative accuracy improvement of critique and recursive critique stages compared to the response stage. Scores are averaged across all datasets. The improvement is calculated as $\exp(\text{Acc}_{\text{stage}}-\text{Acc}_{\text{response}})$, where samples are grouped according to their response accuracy levels.
Nevertheless, we note that these limitations do not diminish the potential of recursive self-critiquing as a scalable oversight paradigm. A lthough current models’ self-critique abilities require improvement, recursive self-critiquing can still yield improvements in weak-to-strong settings as demonstrated in Section 5. This aligns with scalable oversight scenarios where AI provides effective supervision signals when superior to humans.
[…]
Figure 5: AI generartion template in Response Stage
Prompt and hint for $C^{1}$ Generation Input:
[User Prompt]
{question}
[The Start of Response A]
{answer_a}
[The End of Response A]
[The Start of Response B]
{answer_b}
[The End of Response B]
You are given a question and two responses. You should first think step by step and decide which response is better. Avoid any positional bias or length bias and only focus on the quality of the responses. Output your final choice by strictly following this format: "[[A]]" if response A is better. "[[B]]" if response B is better. HINT: Let me carefully analyze which response is better. Firstly, the response
Figure 6: Prompt and hint for $C^{1}$ Generation in AI experiments
Prompt and hint for $C^{2}$ Input:
[User Prompt]
{question}
[The Start of Response A]
{answer_a}
[The End of Response A]
[The Start of Response B]
{answer_b}
[The End of Response B]
[The Start of Critic A]
{critic_a}
[The End of Critic A]
[The Start of Critic B]
{critic_b}
[The End of Critic B]
You are given a question, two responses, and two critics of the responses. You should first think step by step and decide which critics is better. Avoid any positional bias or length bias and only focus on the quality of the critics. Output your final choice by strictly following this format: "[[A]]" if critic A is better. "[[B]]" if critic B is better. HINT: Let me carefully analyze which critic is better. Firstly, the critic
Figure 7: Prompt and hint for $C^{2}$ in AI experiments
Prompt and hint for $C^{3}$ Input:
[User Prompt]
{question}
[The Start of Response A]
{answer_a}
[The End of Response A]
[The Start of Response B]
{answer_b}
[The End of Response B]
[The Start of Critic A]
{critic_a}
[The End of Critic A]
[The Start of Critic B]
{critic_b}
[The End of Critic B]
[The Start of Critic of Critic A]
{critic_of_critic_a}
[The End of Critic of Critic A]
[The Start of Critic of Critic B]
{critic_of_critic_b}
[The End of Critic of Critic B]
You are given a question, two responses, and two critics of the responses, and the two critics of the critics. You should first think step by step and decide which critics of critic is better. Avoid any positional bias or length bias and only focus on the quality of the critics of critic. Output your final choice by strictly following this format: "[[A]]" if critic of critic A is better. "[[B]]" if critic of critic B is better. HINT: Let me carefully analyze which critic of critic is better. Firstly, the critic of critic
Figure 8: Prompt and hint for $C^{3}$ in AI experiments
Prompt for Response Generation Please answer the following multiple-choice question. Your response should include the following sections: - Explanation of Choice: Provide a concise explanation of why this option is chosen, including specific reasons or evidence supporting this choice, starts with ‘Explanation: ‘ within 256 words. - Analysis of Other Options: Analyze each of the remaining options one by one, and explain why they are less suitable than the chosen answer within 256 words. - Answer: On a separate line, starts with ‘Answer: ‘, state your chosen option (A, B, C, or D) only, without any additional text. ### Question: {question} ### Options: {options} Example Input: ### Question: What is the largest continent in the world? ### Options: A. Antarctica B. Africa C. Asia D. South America Example Output: Explanation: Asia is the largest continent in the world by area, covering approximately 44.57 million square kilometers. It is widely recognized in the geographical community as the largest continent. Analysis of Other Options: A) Antarctica: Although Antarctica is very large, it is smaller than Asia and is not usually ranked by land area in this context. B) Africa: Africa is the third-largest continent, but it is smaller than Asia. D) South America: South America is even smaller, making it an incorrect choice for this question. Answer: C
Figure 9: AI generartion template in Response Stage
Prompt for Critic Generation You are given a multiple-choice question and two responses from different individuals. Each response includes the person’s chosen answer and their explanation. Your task is to identify which person’s answer is correct based on their explanations and the information known about the question. Follow this structure for your response: - Explanation of Choice: Compare both explanations to your knowledge about the topic and determine which aligns better with the correct answer, starts with ‘Explanation: ‘. - Analysis of Other Options: Review the explanation provided by each person. Evaluate the reasoning and evidence behind each choice and point out any inaccuracies or correct assumptions. - Answer: On a separate line, starts with ‘Answer: ‘, state your chosen option (A, B, C, or D) only, without any additional text. ### Question: {question} ### Options: {options} ### Person 1’s Response: {gen1} ### Person 2’s Response: {gen2} Example Input: ### Question: Which element has the atomic number 6? ### Options: A) Nitrogen B) Oxygen C) Carbon D) Helium ### Person 1’s Response: Chosen Answer: C Explanation: Carbon is the element with atomic number 6, well-known for being the basis of organic chemistry. ### Person 2’s Response: Chosen Answer: A Explanation: Nitrogen is important for life on Earth, making up a large portion of the atmosphere. Example Output: Explanation: - Person 1 accurately states that Carbon is the element with atomic number 6, supporting their choice with the relevance to organic chemistry. - Person 2 incorrectly chooses Nitrogen, which has an atomic number of 7, misunderstanding the atomic number. Person 1’s explanation aligns correctly with the atomic properties of elements, as Carbon indeed has the atomic number 6. Answer: C
Figure 10: AI generartion template in Critic Stage
Prompt for $C^{2}$ Generation You are given a multiple-choice question. And two individuals, Person 1 and Person 2, have selected their answers and provided their explanations for their choices. Additionally, two more individuals, Reviewer 1 and Reviewer 2, have read these explanations and provided their evaluations of Person 1’s and Person 2’s reasoning. Your task is to identify which answer is correct based on their explanations and the information known about the question. Follow this structure for your response: - Explanation of Choice: Compare both explanations to your knowledge about the topic and determine which aligns better with the correct answer, starts with ‘Explanation: ‘. - Analysis of Other Options: Review the explanation provided by each person. Evaluate the reasoning and evidence behind each choice and point out any inaccuracies or correct assumptions. - Answer: On a separate line, starts with ‘Answer: ‘, state your chosen option (A, B, C, or D) only, without any additional text. ### Question: {question} ### Options: {options} ### Person 1’s Response: {gen1} ### Person 2’s Response: {gen2} ### Reviewer 1’s Response: {c1} ### Reviewer 2’s Response: {c2}
Figure 11: AI generartion template in $C^{2}$ Stage (Part 1)
Prompt for $C^{2}$ Generation Example Input: ### Question: Which element has the atomic number 6? ### Options: A) Nitrogen  B) Oxygen  C) Carbon  D) Helium ### Person 1’s Response: Chosen Answer: C Explanation: Carbon is the element with atomic number 6, well-known for being the basis of organic chemistry. ### Person 2’s Response: Chosen Answer: A Explanation: Nitrogen is important for life on Earth, making up a large portion of the atmosphere. ### Reviewer 1’s Response: Explanation: Person 1 accurately states that Carbon has the atomic number 6. This explanation is correct and directly addresses the question. Carbon is indeed a fundamental element … Answer: C ### Reviewer 2’s Response: Explanation: Person 1 has correctly identified Carbon as the element with atomic number 6. Their … Answer: C Example Output: Explanation: - Reviewer 1 provides a clear and direct evaluation of Person 1’s and Person 2’s responses. They … Answer: C
Figure 12: AI generartion template in $C^{2}$ Stage (Part 2)
Prompt for $C^{3}$ Generation (Part 1) You are given a multiple-choice question, along with responses from two individuals, Person 1 and Person 2, their explanations, evaluations from two reviewers (Reviewer 1 and Reviewer 2), and now evaluations of reviewers from two additional judges (Judger 1 and Judger 2). Your task is to identify which answer is correct, based on judges’ explanations and the final evaluations from the judges. Follow this structure for your response: - Explanation of Choice: Based on the information provided by judges, compare both explanations and determine which one aligns better with the correct answer. Start with ‘Explanation: ‘. Consider both the critiques from Judger 1 and Judger 2. - Analysis of Reviewer Evaluations: Evaluate the critiques provided by Judger 1 and Judger 2. Discuss their reasoning and whether it aligns with the correct answer. Consider whether Judger 1 and Judger 2’s decisions reflect the validity of these critiques. - Answer: On a separate line, start with ‘Answer: ‘, stating your final chosen option (A, B, C, or D) only, without any additional text. ### Question: {question} ### Options: {options} ### Person 1’s Response: {gen1} ### Person 2’s Response: {gen2} ### Reviewer 1’s Response: {c1} ### Reviewer 2’s Response: {c2} ### Judger 1’s Response: {j1} ### Judger 2’s Response: {j2}
Figure 13: AI generation template in $C^{3}$ Stage with Judger Evaluations (Part 1)
Prompt for $C^{3}$ Generation (Part 2) Example Input: ### Question: Which element has the atomic number 6? ### Options: A) Nitrogen  B) Oxygen  C) Carbon  D) Helium ### Person 1’s Response: Chosen Answer: C Explanation: Carbon is the element with atomic number 6, well-known for being the basis of organic chemistry. ### Person 2’s Response: Chosen Answer: A Explanation: Nitrogen is important for life on Earth, making up a large portion of the atmosphere. ### Reviewer 1’s Response: Chosen Answer: C Explanation: Person 1 accurately states that Carbon has the atomic number 6. This explanation is correct and directly addresses the question. Carbon is indeed a fundamental element in organic chemistry. ### Reviewer 2’s Response: Chosen Answer: C Explanation: Person 1 has correctly identified Carbon as the element with atomic number 6. Their explanation is scientifically accurate and directly answers the question. ### Judger 1’s Response: Chosen Answer: C Explanation: Based on Reviewer 1 and Reviewer 2’s critique, Person 1’s explanation is indeed correct. Nitrogen (A) does not have atomic number 6, so Person 2’s response is invalid. I agree with Person 1’s answer. ### Judger 2’s Response: Chosen Answer: C Explanation: After considering Reviewer 2’s feedback and Judger 1’s decision, it is clear that Carbon (C) is the correct answer. Person 1’s explanation holds up against the reviewers’ critique. I agree with Person 1’s answer. Example Output: Explanation: - Both Reviewer 1 and Reviewer 2 agree that Person 1’s explanation is scientifically accurate, and Judger 1 and Judger 2 both reaffirm this conclusion. Based on this consensus, Person 1’s explanation aligns with the correct answer. Answer: C
Figure 14: AI generation template in $C^{3}$ Stage with Judger Evaluations (Part 2)
Table 6: High quality and low quality response examples.
Quality Definition Type Example and Translation
High quality Contains three elements: textual evidence, reasoning, and conclusion. Clear and coherent expression with logical flow. English Origin: 根据题中的"before the end of the century"可定位到原文"Scientists have already pointed out that unless something … before this century is out"。从中可以得知如果不采取措施限制人口快速增长或开发新的食物来源,数百万人将在本世纪结束前死于饥饿。因此可推断作者认为世界最大的问题是如何养活日益增长的人口,选B。 Translated: Based on the phrase "before the end of the century", we can locate "Scientists have already pointed out that unless something … before this century is out". This indicates that without measures to limit population growth or develop new food sources, millions will face starvation. Therefore, feeding the growing population appears to be the major challenge, supporting option B.
Chinese Origin: 文章第三段说:"由于杂交水稻不同熟期组合的出现,全国各地涌现出各种与杂交水稻种植相配套的新型种植模式。"杂交水稻和新型种植模式的出现是因果关系,而不是正好与新型种植模式相配,所以选D。 Translated: The third paragraph states: "Due to the emergence of hybrid rice varieties with different maturity periods, new planting patterns have emerged nationwide to match hybrid rice cultivation." The relationship between hybrid rice and new planting patterns is causal, not just coincidental matching, therefore D is correct.
Math Origin: $f(x)=2\cos^{2}x-\sin^{2}x+2$$\cos 2x=2\cos^{2}x-1$$2\cos^{2}x=\cos 2x+1$$\sin^{2}x+\cos^{2}x=1$$\sin^{2}x=(1-\cos 2x)/2$$f(x)=\frac{3}{2}\cos 2x+\frac{5}{2}$。通过周期计算和最值分析,得到答案B。 Translated: First simplify $f(x)=2\cos^{2}x-\sin^{2}x+2$. Using double angle formula $\cos 2x=2\cos^{2}x-1$, we get $2\cos^{2}x=\cos 2x+1$. Since $\sin^{2}x+\cos^{2}x=1$, we derive $\sin^{2}x=(1-\cos 2x)/2$. Finally $f(x)=\frac{3}{2}\cos 2x+\frac{5}{2}$. Through period calculation and maximum analysis, we arrive at answer B.
Low quality Missing key elements, unclear reasoning, or lack of evidence support. English Origin: 文章第一句"The gift of being able to describe a face accurately is a rare one"就点明文章主要内容为A。 Translated: The first sentence "The gift of being able to describe a face accurately is a rare one" directly points to option A.
Chinese Origin: 答案C错在:那些已经被认定,应…,原文说的是这种代代相传的非物质文化遗产得到创新(过程中),同时使他们自己具有一种认同感和历史感。 Translated: Option C is wrong because: those already recognized should…, the text actually discusses how this inherited intangible cultural heritage achieves innovation while maintaining a sense of identity and history.
[…]
Table 7: High quality and low quality critique annotation examples.
Question Response1 Response2
根据原文提供的信息,下列选项中,没有体现传统技艺传承价值的一项是 A.农业生产的犁、耙、锄等工具的利用 B.日常生活中医药的陶罐、竹器等物品的制作 C.社会行业中的造纸、印刷、染织等手工艺 D.民俗文化中的剪纸、皮影、年画等艺术的创作 作者主要介绍了传统技艺作为非物质文化遗产有着丰富的文化内涵和艺术价值,传统技艺中具有的民生价值、艺术价值、历史文化价值均是其传承价值所在。 选项A提到的内容没有体现传统技艺传承价值,原因是文章提虽然到了犁、耙、锄等工具的制作是传统手工艺的印记,体现了传统技艺传承价值,但选项重点在于对工具的利用,这并未体现。相关的内容是[农业耕作所需要的犁、耙、锄,饮食起居所需要的陶罐、簸箕、箩筐,交通运输所需要的担、船、轿,以及织锦、染布、造纸等日常生活所需要的各种物品,都留下了手工艺的印记]。因此,我认为答案是「A」。 文章主要探讨了传统技艺的定义,价值和重要性 选项A可能正确的原因:文章中提到「农业耕作所需要的犁、耙、锄」,「具有民生价值」相关的内容说明这些农业生产的相关工具可以体现传统技艺传承价值。此外,选项C可以由原文中「以及织锦、染布、造纸等日常生活所需要的各种物品」,「具有民生价值、艺术价值」的相关内容得作证,而选项D可以由「木雕、石雕、玉雕、泥塑、面塑给人以美的愉悦,剪纸、刻纸、皮影、年画有丰富的文化内涵和艺术价值」,「具有艺术价值、历史文化价值」得到佐证。 而选项B是错误的原因可参考原文「饮食起居所需要的陶罐、簸箕、箩筐」,可以看到原文是饮食起居的陶罐,而不是医药的陶罐。 综合上述分析,答案是「B」。
High Quality Examples
不同意Response1的答案(A)。Response1引用了相关的段落,但认为相关内容「重点在于对工具的利用」,而不是体现传统技艺传承价值。但他引用的内容里提到了「留下了手工艺的印记」,我认为这个内容正体现了农业生产的犁、耙、锄是体现传统技艺传承价值的。认可Response 2的答案(B)。 Response2分析了选项B,并指出「原文提到的是饮食起居的陶罐,而不是医药的陶罐」,他引用了文章中对应内容确实和选项是不符合的。此外,他还分析了选项ACD,并找到了对应的参考内容佐证。但他对于选项A的分析需要纠正,他引用的文章内容「农业生产的犁、耙、锄等工具的利用」和「具有的民生价值」并不在一个段落内,不能论证他的分析。但第四段的开头「手工艺始于对自然物质的利用和各种工具的制造和使用」之后出现的相关描述确实说明了选择A体现了传统技艺传承价值。因此答案应该是(B)。
认可Response 2的答案(B),他引用的文章内容「饮食起居所需要的陶罐、簸箕、箩筐」和选项B「医药的陶罐、竹器等物品」确实是不一致的。而Response1的选择(A),他引用[农业耕作所需要的犁、耙、锄,饮食起居所需要的陶罐、簸箕、箩筐,…都留下了手工艺的印记]里的"手工艺"我认为是传统技艺的一部分,因此不应该选答案A。因此答案应该是(B)。
认可Response 2的答案(B),他引用了文章内容「饮食起居所需要的陶罐、簸箕、箩筐」,而选项B「医药的陶罐、竹器」明显是不一致的,这很可能是题目设下的陷阱,因此Response的2答案应该是对的。而Response1虽然对选项(A)可能是错误的原因做了分析,但基于前面对答案B的分析,它明显是对,因此不该选选项A。因此答案应该是(B)。
Low Quality Examples
认可response 1的答案,不同意Response2的答案,答案应该是(B)。
Table 8: English translations of high quality and low quality critique annotation examples in Table 7.
Question Response1 Response2
According to the text, which of the following does NOT reflect the inheritance value of traditional craftsmanship? A. The use of agricultural tools like plows, harrows, and hoes B. The making of medical pottery and bamboo utensils in daily life C. The handicrafts in social industries such as papermaking, printing, and dyeing D. The folk art creations such as paper-cutting, shadow puppets, and New Year paintings The author mainly introduces traditional crafts as intangible cultural heritage with rich cultural connotations and artistic value. Traditional crafts possess livelihood value, artistic value, and historical-cultural value as their inheritance value. Option A does not reflect traditional craft inheritance value. While the text mentions the tools like plows, harrows, and hoes as traditional handicraft marks, reflecting craft inheritance value, the option focuses on tool usage, which is not demonstrated. The relevant content states [Agricultural tools like plows, harrows, and hoes, kitchenware like pottery, baskets, and hampers, transportation tools like carriers, boats, and sedan chairs, and daily necessities like brocade, dyed cloth, and paper, all bear handicraft marks]. Therefore, I believe the answer is "A". The article mainly discusses the definition, value, and importance of traditional crafts. For Option A: The text mentions "agricultural tools like plows, harrows, and hoes" and their "livelihood value", suggesting these tools reflect traditional craft inheritance value. Additionally, Option C is supported by "daily necessities like brocade, dyed cloth, and paper" showing "livelihood and artistic value", while Option D is evidenced by "wood carving, stone carving, jade carving, clay sculpture, and dough modeling bring aesthetic pleasure, while paper-cutting, paper carving, shadow puppets, and New Year paintings have rich cultural connotations and artistic value". Option B is incorrect as the text mentions "kitchenware pottery" rather than medical pottery. Based on this analysis, the answer is "B".
High Quality Examples
Disagree with Response 1’s answer (A). While Response 1 cites relevant passages, its interpretation that "focus is on tool usage" misses the point about traditional craft inheritance value. The cited phrase "left craftmanship marks" actually demonstrates that agricultural tools reflect traditional craft value. Agree with Response 2’s answer (B). Response 2 correctly analyzes Option B, noting that the text mentions "kitchenware pottery" rather than "medical pottery", with accurate textual evidence. They also provide well-supported analysis for Options A, C, and D. However, their reasoning for Option A needs correction - the connection between "agricultural tools" and "livelihood value" isn’t supported by being in different paragraphs. Nevertheless, the fourth paragraph’s opening about "handicrafts beginning with the use of natural materials and tool manufacturing" supports that Option A reflects traditional craft value. Therefore, the answer should be (B).
Agree with Response 2’s answer (B). Their citation of "kitchenware pottery" from the text clearly contradicts Option B’s "medical pottery". Regarding Response 1’s choice of (A), the reference to "handicraft marks" in the passage about agricultural tools suggests this is part of traditional craftsmanship, so A should not be the answer. Therefore, the answer is (B).
Support Response 2’s answer (B). They correctly identified the discrepancy between "kitchenware pottery" in the text and "medical pottery" in Option B, which appears to be an intentional distinction. While Response 1 analyzes why Option A might be incorrect, based on the analysis of Option B being clearly wrong, Option A cannot be the answer. Therefore, the answer should be (B).
Low Quality Examples
Agree with Response 1, disagree with Response 2, the answer should be (B).
Table 9: High quality and low quality higher-order critic annotation examples.
Question Response 1 Response 2
根据原文提供的信息,下列选项中,没有体现传统技艺传承价值的一项是 A.农业生产的犁、耙、锄等工具的利用 B.日常生活中医药的陶罐、竹器等物品的制作 C.社会行业中的造纸、印刷、染织等手工艺 D.民俗文化中的剪纸、皮影、年画等艺术的创作 作者主要介绍了传统技艺作为非物质文化遗产有着丰富的文化内涵和艺术价值,传统技艺中具有的民生价值、艺术价值、历史文化价值均是其传承价值所在。 选项A提到的内容没有体现传统技艺传承价值,原因是文章提虽然到了犁、耙、锄等工具的制作是传统手工艺的印记,体现了传统技艺传承价值,但选项重点在于对工具的利用,这并未体现。相关的内容是[农业耕作所需要的犁、耙、锄,饮食起居所需要的陶罐、簸箕、箩筐,交通运输所需要的担、船、轿,以及织锦、染布、造纸等日常生活所需要的各种物品,都留下了手工艺的印记]。因此,我认为答案是「A」。 文章主要探讨了传统技艺的定义,价值和重要性 选项A可能正确的原因:文章中提到「农业耕作所需要的犁、耙、锄」,「具有民生价值」相关的内容说明这些农业生产的相关工具可以体现传统技艺传承价值。此外,选项C可以由原文中「以及织锦、染布、造纸等日常生活所需要的各种物品」,「具有民生价值、艺术价值」的相关内容得作证,而选项D可以由「木雕、石雕、玉雕、泥塑、面塑给人以美的愉悦,剪纸、刻纸、皮影、年画有丰富的文化内涵和艺术价值」,「具有艺术价值、历史文化价值」得到佐证。 而选项B是错误的原因可参考原文「饮食起居所需要的陶罐、簸箕、箩筐」,可以看到原文是饮食起居的陶罐,而不是医药的陶罐。 综合上述分析,答案是「B」。
Critic 1 Critic 2
不同意Response1的答案(A)。Response1引用了相关的段落,但认为相关内容「重点在于对工具的利用」,而不是体现传统技艺传承价值。但他引用的内容里提到了「留下了手工艺的印记」,我认为这个内容正体现了农业生产的犁、耙、锄是体现传统技艺传承价值的。 Response2分析了选项B,并指出「原文提到的是饮食起居的陶罐,而不是医药的陶罐」,他引用了文章中对应内容确实和选项是不符合的。此外,他还分析了选项ACD,并找到了对应的参考内容佐证。因此答案应该是(B)。 认可Response 2的答案(B),他引用的文章内容「饮食起居所需要的陶罐、簸箕、箩筐」和选项B「医药的陶罐、竹器等物品」确实是不一致的。而Response1的选择(A),他引用[农业耕作所需要的犁、耙、锄,饮食起居所需要的陶罐、簸箕、箩筐,…都留下了手工艺的印记]里的"手工艺"我认为是传统技艺的一部分,因此不应该选答案A。因此答案应该是(B)。
High Quality Examples
认可Critic 1和2的答案(B),两个Critc都指出答案是B的原因是:文章内容「饮食起居所需要的陶罐、簸箕、箩筐」和选项B「医药的陶罐、竹器等物品」的不一致,因此没有体现传统技艺传承价值。
认可Critic 1和2关于答案(B)的分析,文章内容「饮食起居所需要的陶罐、簸箕、箩筐」和选项B「医药的陶罐、竹器等物品」不一致。但Critic2对于Response1对于选项A错误之处的分析,我觉得理由不充分,「手工艺的印记]不一定直接和「传统技艺」关联,但主要下判断的原因是选项B明显是正确答案。
Low Quality Examples
Critc 1/2的答案是对,应该是(B)。
Table 10: English translations of high quality and low quality higher-order critic annotation examples in Table 9.
Question Response 1 Response 2
According to the text, which of the following does NOT reflect the inheritance value of traditional craftsmanship? A. The use of agricultural tools like plows, harrows, and hoes B. The making of medical pottery and bamboo utensils in daily life C. The handicrafts in social industries such as papermaking, printing, and dyeing D. The folk art creations such as paper-cutting, shadow puppets, and New Year paintings The author mainly introduces traditional crafts as intangible cultural heritage with rich cultural connotations and artistic value. Traditional crafts possess livelihood value, artistic value, and historical-cultural value as their inheritance value. Option A does not reflect traditional craft inheritance value. While the text mentions tools like plows, harrows, and hoes as traditional handicraft marks, reflecting craft inheritance value, the option focuses on tool usage, which is not demonstrated. The relevant content states [Agricultural tools like plows, harrows, and hoes, kitchenware like pottery, baskets, and hampers, transportation tools like carriers, boats, and sedan chairs, and daily necessities like brocade, dyed cloth, and paper, all bear handicraft marks]. Therefore, I believe the answer is "A". The article mainly discusses the definition, value, and importance of traditional crafts. For Option A: The text mentions "agricultural tools like plows, harrows, and hoes" and their "livelihood value", suggesting these tools reflect traditional craft inheritance value. Additionally, Option C is supported by "daily necessities like brocade, dyed cloth, and paper" showing "livelihood and artistic value", while Option D is evidenced by "wood carving, stone carving, jade carving, clay sculpture, and dough modeling bring aesthetic pleasure, while paper-cutting, paper carving, shadow puppets, and New Year paintings have rich cultural connotations and artistic value". Option B is incorrect as the text mentions "kitchenware pottery" rather than medical pottery. Based on this analysis, the answer is "B".
Critic 1 Critic 2
Disagree with Response 1’s answer (A). While Response 1 cites relevant passages, its interpretation that "focus is on tool usage" misses the point about traditional craft inheritance value. The cited phrase "left craftmanship marks" actually demonstrates that agricultural tools reflect traditional craft value. Response 2 correctly analyzes Option B, noting that the text mentions "kitchenware pottery" rather than "medical pottery", with accurate textual evidence. They also provide well-supported analysis for Options A, C, and D. Therefore, the answer should be (B). Agree with Response 2’s answer (B). Their citation of "kitchenware pottery" from the text clearly contradicts Option B’s "medical pottery". Regarding Response 1’s choice of (A), the reference to "handicraft marks" in the passage about agricultural tools suggests this is part of traditional craftsmanship, so A should not be the answer. Therefore, the answer is (B).
High Quality Examples
Agree with both Critics’ answer (B). Both critics point out that the discrepancy between "kitchenware pottery" in the text and "medical pottery" in Option B shows it does not reflect traditional craft inheritance value.
Agree with both Critics’ analysis of option B, noting the clear difference between "kitchenware pottery" in the text and "medical pottery" in the option. However, Critic 2’s reasoning about Response 1’s option A analysis is insufficient - "handicraft marks" doesn’t necessarily equate to "traditional crafts", though this doesn’t affect the final judgment as option B is clearly correct.
Low Quality Examples
Critics 1/2 are correct, the answer should be (B).